# Human limits in Machine Learning: Prediction of plant phenotypes using soil microbiome data

### v1

@@ B.2.4 Robust prediction in imbalanced datasets with data augmentation
For the FMS decision tree for the RF model (Figure 6), the highest weighted F1 score (0.934 with 0.5% of the data) is achieved with data augmentation, normalization/zero replacement strategy #6 (CSS+pseudo), and Order level. Another path of the decision tree follows data augmentation and any normalization/zero replacement strategy except #6 (CSS+pseudo), #14 (rarefy+pseudo), and #18 (clr+pseudo) which yields an average weighted F1 scores of 0.892 for 42.5% of the data preprocessing options. If data augmentation is not an option (left of the root), the highest weighted F1 score available is 0.868 with Phylum level, and any normalization/zero replacement strategy except #18 (clr+pseudo) or #20 (clr+bayesMult). For the FMS decision trees on the other responses, see Figures S16 to S20 in the Appendix.
@@
Similarly, in Figure 7 for the BNN model, the highest weighted F1 score (0.896 with 15% of the data preprocessing options) is achieved when we do data augmentation, we use any taxonomic level except Phylum, and we use any normalization/zero replacement strategy except #10 (COM+pseudo) and #18 (clr+pseudo). See Figures S21 to S25 in the Appendix for other responses.
@@
While the specific recommendations on normalization, zero replacement and taxonomic level are model-specific, both models perform better with data augmentation. In terms of taxonomic level, we note that the BNN was only run on Phylum and Class, and thus, the highest accuracy is obtained with Class level (when Phylum$=0$ is true). This does not contradict the result from the RF that identified Order level as the one yielding higher accuracy. We cannot rule out that the BNN would also have higher accuracy with Order compared to Class. The results from the RF, though, seem to suggest that there is a peak at Order, and more granularity in Family and Genus does not seem to provide more predictive power.
@@
Figure 6: Full model selection decision tree with a maximum depth of 4 summarizing the results of random forest models on pitted scab disease (Scabpit). When the condition at a node is true, we follow the branch on the left, and when the condition is false, we follow the branch on the right. The percentage of the data preprocessing options and mean of weighted F1 scores are shown in each nodes. The darker the red, the higher the weighted F1 score.
B.3 Prediction of Potato Disease from Environmental Data
One of the questions to address in our work is whether prediction accuracy is improved by the inclusion of microbiome data, or if environmental factors provide enough signal to classify potatoes in diseased or non-diseased groups.
@@ B.3 Prediction of Potato Disease from Environmental Data
We found that environmental factors indeed provide sufficient signal to predict pitted scab disease as illustrated in Figure 8 which shows the weighted F1 scores by RF and BNN models based on environmental (soil characteristics) data for pitted scab. The range of each boxplot corresponds to the six scaling methods described in Section D.1.2. In contrast with the normalization methods in microbiome data, we observe here that the scaling methods do not seem to have an effect on prediction as evidenced by narrow boxplots, and that weighted F1 scores are all higher than 0.7, and therefore, comparable to the models fitted on microbiome data alone. These results suggest that environmental factors alone are powerful to predict the incidence of pitted scab in the tubers. As microbiome data is more expensive than environmental data, we suggest to prefer environmental predictors under restricted monetary budget. See Figures S26 to S30 in the Appendix for other responses.
Figure 8: Boxplots of the weighted F1 scores by random forest (RF) and Bayesian neural network (BNN) models for pitted scab disease (Scabpit) using environmental predictors.
B.4 Leveraging Microbiome and Environmental Data in the Prediction of Potato Disease
@@
As expected, prediction accuracy is improved by the inclusion of both datatypes: microbiome and environmental. Figure 9 shows the weighted F1 scores by RF and BNN models based on combined datasets with environmental and microbial predictors for pitted scab. We only focus on the most accurate models identified in Section B.1. First, we note that a model that uses OTU abundances outperforms a model that uses alpha diversity as a predictor (comparison of Alpha with OTU-S3) for both types of models (RF and BNN). This suggests that we lose information by transforming abundances into diversity measures. Second, models including only OTU abundances (OTU-S3) perform comparably to models that include both types of predictors (OTU-S3+Soil+DS) which suggests that the microbial data indeed has substantial predictive power on its own, but adding microbiome to soil predictors may not provide much benefit for high predictive power, with the only exception of Phylum level OTU-S3+Soil+DS in a RF model (Figure 9). Generally, the model with only soil information (shown as a blue dashed line) performs just as accurately. Third, contrary to prior expectations that microbial communities at finer resolution would be a better choice for predicting pitted scab or other diseases, our study does not find any evidence that the prediction power increases when moving up from Phylum to Genus level. Particularly, the prediction power of OTU-S3 in RF model increases from Class to Genus, and this pattern is not preserved when diversity is used instead of OTU abundances. For example, a model with only alpha diversity as the predictor (Alpha) shows decreasing weighted F1 score as we move from Phylum to Genus level. Both models (RF and BNN) when including all types of predictors (OTU-S3+Soil+DS) result in the similar weighted F1 score regardless of taxonomic level. See Figures S31 to S35 in the Appendix for other responses.
@@
Figure 9: Boxplots with the weighted F1 scores (y-axis) by random forest and Bayesian neural network (Bayesian NN) models (x-axis) for pitted scab disease (Scabpit). The range of each box plot depicts the weighted F1 scores for normalized datasets at each taxonomic level. There are 6 normalization methods for alpha diversity and environmental factors (Soil and DS), and one normalization method for microbiome (OTU-S3). See Table 3 and 4 for description of models and number of normalization methods for each predictors. The models including both types of predictors outperform other models, yet models including microbiome data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, microbiome data is more expensive to collect, and perhaps not necessary, given that the model without microbiome data (Soil) performs just as accurately (blue dashed line).
Appendix C Discussion
@@
Soil microbiome represents the most complex and least understood aspect of soil health. In this study, we use ML techniques such as random forest (RF) and Bayesian neural network (BNN) to determine whether soil microbial information has any predictive power for plant outcomes such as yield and disease. Our results show that the microbiome plays a role in predicting potato disease, but the most accurate prediction originates from a combination of both datatypes: microbiome and environmental factors. Given that prediction with environmental factors alone was sufficiently powerful, it is uncertain whether the extra expense to sequence microbiome data is worth the cost.
@@
The prediction power of microbiome data varies depending the outcome we want to predict. For example, among all models that predict diseases, models for the pitted scab disease receive very high weighted F1 scores compared to other diseases. We further confirm this predictive power by comparing the performance of models trained on the real microbiome data to models trained on randomly generated data. Given that the prediction of the pitted scab disease is far from random, we can confidently conclude that this disease can be accurately predicted from microbiome data. It is noteworthy, however, that pitted scab disease is precisely one of the diseases that are easier to be visually detected, and thus, there is a reliable separation among the two classes (diseased and non-diseased) which is aiding in prediction by machine-learning models. Other diseases, and more so yield, do not have such clear distinction between classes which results in lower predictive power. That is, we believe that the lack of prediction accuracy in yield is not driven by a lack of a biological connection between soil microbiome and yield, but on the lack of accurate labels that distinguish the two classes (e.g. low and high yield). If humans cannot distinguish what is low yield vs high yield, then that ambivalence will propagate into the machine-learning classification. This conclusion seems to be confirmed when we notice that yield cannot be accurately predicted by any of 14 models in consideration (Table 3). We conclude that one of the main challenges when applying ML methods in biological applications is the artificial binarization of phenotypes. More work is needed to improve the performance of regression models that can predict continuous phenotypes when faced with limited sample sizes that are common in biological domains.
@@
We show that the normalization and zero replacement strategies for microbiome data sets have a huge impact on results, and unfortunately, we could not find any pattern to suggest that one strategy is superior to others. The full model selection (FMS) decision tree, however, suggests that certain normalization methods such as centered log-ratio normalization with pseudo-zero replacement (clr+pseudo) or centered log-ratio normalization with Bayesian-multiplicative treatment (clr+bayesMult) should not be used for disease prediction for future studies. We recommend domain scientists to perform different types of normalization and zero replacement strategies in their data prior to be used as input for prediction. See also [4, 22, 7] for other performance comparisons on types of normalizations.
@@
In terms of feature selection, we considered different strategies to select important features (ML, network comparison, and intersection of both). There is significant overlap between important OTUs by two methods and the inclusion of this subset of predictors allowed us to build less complex models with comparable good performance (Figure 4 for pitted scab disease). We did not observe one feature selection strategy that outperformed the others. For example, the weighted F1 scores obtained when using the OTUs identified as important by the ML methods are comparable to the weighted F1 scores when including OTUs identified as important by network comparison. It is worth mentioning, however, that the performance is also comparable to that obtained when including all OTUs which shows that the feature selection strategies work at preserving the OTUs that have predictive signal while simultaneously allowing computationally expensive models (like BNN) to be applied. Tables S2 and S3 in the Appendix contain the important OTUs identified by the two feature selection methods for the Phylum level associated with pitted scab disease.
@@
In our study, we could not identify an overall prediction superiority on certain taxonomic levels compared to others. While it is intuitive to expect that finer taxonomic levels would provide more power to predict potato outcomes, this is only true in some cases such as the OTU-S3 model in RF and the OTU-S1 model in BNN (Figure 4) where Genus level indeed outperforms the other taxonomic levels. However, we cannot find evidence to support this expectation in all other cases.
Even though this finding disagrees with common expectations in microbiome research, we believe that it is a positive result from the machine learning perspective. If less refined taxonomic levels perform just as well as finer taxonomic levels, then the model could be fitted with fewer features, and would thus be less computational expensive and more efficient. Furthermore, the more features in the data, the more samples are needed to train the models. So, ML models that are trained on limited samples sizes (as the ones we have in this study) would benefit greatly by having fewer features.
@@
When including environmental features such as soil physicochemical properties and microbial population density of soil in the model, we achieved higher weighted F1 score values. In general, we investigated 14 different models for yield and disease prediction with different combinations of microbiome and environmental data. Results show poor performance in predicting yield across different models. The best results for pitted scab are achieved by combining alpha diversity and soil chemistry (Alpha+Soil) for RF and important OTUs and soil data (OTU-S3+Soil) for BNN. Although the best-performing models include microbiome predictors (Alpha and OTU-S3), it is important to note that the models without microbiome data are comparably powerful as those including microbiome data. While the models trained with microbiome data alone show that microbiome can effectively predict pitted scab disease, the fact that models without this type of data continue to perform well provides evidence that microbiome data may not be necessary to achieve reasonable prediction. In fact, when collection of microbiome data requires much higher cost investment, the increased prediction accuracy by including microbiome predictors may not be enough to justify the extra cost.
@@
The RF and BNN methods offer similar performance in all models, which validates the results of both methods. However, in this study, BNN take far longer to train due to weight sampling and approximation, especially on more refined taxonomic level such as Order, Family, and Genus. Thus, we believe that based on the current dataset, RF is the preferred decision-making model with high prediction potential for disease outcomes when including a combination of microbiome and environmental predictors.
Appendix D Methods
@@
Figure 1 shows a graphical representation of our pipeline with three major steps: (1) data preparation, (2) feature selection, and (3) classification based on random forest (RF) and Bayesian neural network (BNN) models with various types of predictors, including microbiome data (OTUs), environmental, and a combination of both. We describe each step in the pipeline in the next subsections.
D.1 Data description
@@
In this study, we focus on the soil microbiome (matrix of abundances) in a variety of taxonomic orders, including Phylum, Class, Order, Family, and Genus as well as other environmental information from soil samples acquired from potato fields in Wisconsin and Minnesota. The dataset consists of measurements related to soil health, potato yield and soil quality information. The soil health data are collected from pre-plant fields and include soil physicochemical properties, soil microbiome composition, soil microbiome diversity, and soil pathogen abundance. The potato yield and quality assessments are taken from the following growing season with indicators including tuber yield and disease severity. Overall, we have 256 samples, 108 of which are taken from fields in Minnesota, and 148 of which are taken from fields in Wisconsin. We list all measurements in Table S1 in the Appendix.
D.1.1 Binarization of Response Variables
@@
All six responses in the dataset are continuous, so we need to binarize them to fit the classification models. For the disease-related responses, we simply make the binary label 00 if there is no presence of disease, and $1$ if there was detection of disease (that is, if the continuous response is greater than $0.0$). Binarizing the yield responses variables is harder as there is not a universal standard to classify potato yield to be low or high. Furthermore, yield values are highly dependent on the type of potato variety. We assign a label of 00 (low yield) to samples with a yield less than the variety-specific median. Similarly, we assign a label of $1$ (high yield) to samples with yield greater than the variety-specific median. We illustrate this approach in Figure S1 in the Appendix.
@@
D.1.2 Data Filtering, Normalization, and Zero Replacement
The input data is a matrix with non-negative read counts that were generated by a sequencing procedure. Let $w^{(k)}=[w_{1}^{(k)},...,w_{p}^{(k)}]$ be the total read counts of sample $k$ containing $p$ OTUs, where $w^{(k)}$ is a composition that adds up to a fixed value of $m^{(k)}=\sum_{i=1}^{p}w_{i}^{(k)}$. This value $m^{(k)}$ is the sequencing depth, which varies across samples and is predetermined by technical factors resulting in highly sparse data. It is reasonable to filter out a certain set of OTUs as first data preparation step. For filtering, we only include OTUs that appear in at least 15 samples.
Table 2 displays the number of features (OTUs) before and after filtering for different taxonomic levels.
Level # OTUs # OTUs after filtering
Phylum 57 42
Class 152 108
Order 378 224
Family 476 255
Genus 1319 485
@@ D.1.2 Data Filtering, Normalization, and Zero Replacement
With four zero replacement methods and five normalization methods, we create 20 datasets by the combination of zero replacement and normalization techniques. Our goal is to study the effect of the zero replacement and normalization choice in the performance of the deep learning methods. Namely, we have the following 20 combinations, $\text{NM}_{1}$: TSS+none, $\text{NM}_{2}$: TSS+pseudo, $\text{NM}_{3}$: TSS+multRepl, $\text{NM}_{4}$: TSS+bayesMult, $\text{NM}_{5}$: CSS+none, $\text{NM}_{6}$: CSS+pseudo, $\text{NM}_{7}$: CSS+multRepl, $\text{NM}_{8}$: CSS+bayesMult, $\text{NM}_{9}$: COM+none, $\text{NM}_{10}$: COM+pseudo, $\text{NM}_{11}$: COM+multRepl, $\text{NM}_{12}$: COM+bayesMult, $\text{NM}_{13}$: rarefy+none, $\text{NM}_{14}$: rarefy+pseudo, $\text{NM}_{15}$: rarefy+multRepl, $\text{NM}_{16}$: rarefy+bayesMult, $\text{NM}_{17}$: clr+none, $\text{NM}_{18}$: clr+pseudo,, $\text{NM}_{19}$: clr+multRepl, and $\text{NM}_{20}$: clr+bayesMult. For convenience, we use the notation $\text{NM}_{i}$ (Normalization Method) for $i=1,\dots,20$ in the Full Model Selection section (Section B.2.5).
For the environmental predictors of soil chemistry and microbial population density in the soil, we apply six scaling methods: (1) standardize features by subtracting the mean and scaling to unit variance [8]; (2) scale each feature to a $[0,1]$ range; (3) scale each feature by its maximum absolute value; (4) scale features by subtracting the median and scaling to the interquantile range [32]; (5) transform the features to follow a normal distribution [21]; (6) normalize samples individually to the unit norm.
After normalization, the datasets are split into training, validation, and testing sets with 10-fold cross-validation. We used 80% of samples for training and validation, and 20% for testing.
D.1.3 Data Augmentation
@@
Regarding the data augmentation procedure, instead of simply adding a randomly generated noise to the original sample, we subset the data by variety and label, compute the mean (and standard deviation) abundance value for this subset, and create a new sample that includes the original data plus a Gaussian error with mean $\mu/100$ and standard deviation $\sigma/100$ where $\mu,\sigma$ are the subset-specific mean and standard deviation, respectively. This approach is illustrated in Figure S2 in the Appendix. By the end of this procedure we would have a balanced augmented training set with 400 samples per label for each of the five taxonomic levels (Phylum, Class, Order, Family, and Genus).
D.2 Feature Selection
Feature selection involves the identification of important features (or covariates) that have high predictive power. Given the high-dimensionality of the data (e.g. 256 original samples for 485 OTUs at the Genus level), feature selection is necessary, especially for BNN models that are computationally expensive.
We pursue two approaches for feature selection: (1) using machine-learning models to assess variable importance, and (2) using network analyses. To focus exclusively on the effect of feature selection, we only consider one type of normalization and zero replacement strategy in this investigation, namely, total sum scaling normalization without zero replacement ($\text{NM}_{1}$: TSS+none).
D.2.1 Using ML Models for Feature Selection
@@
After running all seven strategies, we assign a value ("TOTAL") to each OTU based on the number of times the OTU is selected as an important feature under the seven criteria. That is, an OTU that is selected as important by all seven strategies will have a value of 7. The OTUs are sorted based on "TOTAL" column and top 30% of them are selected as important features. Thus, 30, 36, 75, 85, and 162 OTUs are selected for Phylum, Class, Order, Family and Genus levels, respectively. For example, Table S2 in the Appendix shows the top 30 OTUs at Phylum level and by which strategies they are identified as important features to predict the pitted scab response.
D.2.2 Using Network Comparison for Feature Selection

### v2

@@ Results
For the FMS decision tree for the RF model (Figure 6), the highest weighted F1 score (0.934 with 0.5% of the data) is achieved with data augmentation, normalization/zero replacement strategy #6 (CSS+pseudo), and Order level. Another path of the decision tree follows data augmentation and any normalization/zero replacement strategy except #6 (CSS+pseudo), #14 (rarefy+pseudo), and #18 (clr+pseudo) which yields an average weighted F1 scores of 0.892 for 42.5% of the data preprocessing options. If data augmentation is not an option (left of the root), the highest weighted F1 score available is 0.868 with Phylum level, and any normalization/zero replacement strategy except #18 (clr+pseudo) or #20 (clr+bayesMult). For the FMS decision trees on the other responses, see Figures S20 to S24 in the Supplementary Material.
@@
Similarly, in Figure S19 for the Bayesian NN model, the highest weighted F1 score (0.896 with 15% of the data preprocessing options) is achieved when we do data augmentation, we use any taxonomic level except Phylum, and we use any normalization/zero replacement strategy except #10 (COM+pseudo) and #18 (clr+pseudo). See Figures S25 to S29 in the Supplementary Material for other responses.
@@
While the specific recommendations on normalization, zero replacement and taxonomic level are model-specific, both models perform better with data augmentation. In terms of taxonomic level, we note that the Bayesian NN was only run on Phylum and Class, and thus, the highest accuracy is obtained with Class level (when Phylum$=0$ is true). This does not contradict the result from the RF that identified Order level as the one yielding higher accuracy. We cannot rule out that the Bayesian NN would also have higher accuracy with Order compared to Class. The results from the RF, though, seem to suggest that there is a peak at Order, and more granularity in Family and Genus does not seem to provide more predictive power.
@@
Figure 6: Full model selection decision tree with a maximum depth of 4 summarizing the results of random forest models on pitted scab disease. When the condition at a node is true, we follow the branch on the left, and when the condition is false, we follow the branch on the right. The percentage of the data preprocessing options and mean of weighted F1 scores are shown in each node.
Prediction of Potato Disease from Environmental Data
@@
One of the questions to address in our work is whether prediction accuracy is improved by the inclusion of microbiome data, or if environmental factors (usually cheaper to collect) provide enough signal to classify potatoes in diseased or non-diseased groups. We found that environmental factors indeed provide sufficient signals to predict pitted scab disease as illustrated in Figure S30 which shows the weighted F1 scores by RF and Bayesian NN models based on environmental (soil characteristics) data for pitted scab. The range of each boxplot corresponds to the six scaling methods described in Section ‘Data Filtering, Normalization, and Zero Replacement’. In contrast with the normalization methods in microbiome data, we observe here that the scaling methods do not seem to have an effect on prediction as evidenced by narrow boxplots, and that weighted F1 scores are all higher than 0.75, and therefore, comparable to the models fitted on microbiome data alone. These results suggest that environmental factors alone are powerful to predict the incidence of pitted scab in the tubers. As microbiome data is more expensive than environmental data, we suggest to prefer environmental predictors under restricted monetary budget. See Figures S31 to S35 in the Supplementary Material for other responses.
Leveraging Microbiome and Environmental Data in the Prediction of Potato Disease
@@
As expected, prediction accuracy is improved by the inclusion of both datatypes: microbiome and environmental. Figure 7 shows the weighted F1 scores by RF and Bayesian NN models based on combined datasets with environmental and microbial predictors for pitted scab. We only focus on the most accurate models identified in Section ‘Overall performance of predictive models: Manual binarization causes inaccurate prediction of yield’. First, we note that a model that uses OTU abundances outperforms a model that uses alpha diversity as a predictor (comparison of Alpha with OTU-S3) for both types of models (RF and Bayesian NN). This suggests that we lose information by transforming abundances into diversity measures. Second, models including only OTU abundances (OTU-S3) perform comparably to models that include both types of predictors (OTU-S3+Soil+DS) which suggests that the microbial data indeed has substantial predictive power on its own, but adding microbiome to soil predictors may not provide much benefit for high predictive power, with the only exception of Phylum level OTU-S3+Soil+DS in a RF model (Figure 7). Generally, the model with only soil information (shown as a blue dashed line) performs just as accurately. Third, contrary to prior expectations that microbial communities at finer resolution would be a better choice for predicting pitted scab or other diseases, our study does not find any evidence that the prediction power increases when moving up from Phylum to Genus level. Particularly, the prediction power of OTU-S3 in RF model increases from Class to Genus, and this pattern is not preserved when diversity is used instead of OTU abundances. For example, a model with only alpha diversity as the predictor (Alpha) shows decreasing weighted F1 score as we move from Phylum to Genus level. Both models (RF and Bayesian NN) when including all types of predictors (OTU-S3+Soil+DS) result in the similar weighted F1 score regardless of taxonomic level. See Figures S36 to S40 in the Supplementary Material for other responses.
@@
Figure 7: Boxplots with the weighted F1 scores (y-axis) by random forest and Bayesian neural network (Bayesian NN) models for pitted scab disease. The range of each box plot depicts the weighted F1 scores for normalized datasets at each taxonomic level. There are 6 normalization methods for alpha diversity and environmental factors (Soil and DS), and one normalization method for microbiome (OTU-S3). See Table 2 and 3 for description of models and number of normalization methods for each predictors. The models including both types of predictors outperform other models, yet models including microbiome data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, microbiome data is more expensive to collect, and perhaps not necessary, given that the model without microbiome data (Soil) performs just as accurately (blue dashed line).
Discussion
@@
Soil microbiome represents the most complex and least understood aspect of soil health. In this study, we use ML techniques such as random forest (RF) and Bayesian neural network (Bayesian NN) to determine whether soil microbial information has any predictive power for plant outcomes such as yield and disease. Our results show that the microbiome plays a role in predicting potato disease, but the most accurate prediction originates from a combination of both datatypes: microbiome and environmental factors. Given that prediction with environmental factors alone was sufficiently powerful, it is uncertain whether the extra expense to sequence microbiome data is worth the cost.
@@
The prediction power of microbiome data varies depending on the outcome we want to predict. For example, among all models that predict diseases, models for the pitted scab disease receive very high weighted F1 scores compared to other diseases. We further confirm this predictive power by comparing the performance of models trained on the real microbiome data to models trained on randomly generated data (see Figures 7 and 8 in the Supplementary Material). Given that the prediction of the pitted scab disease is far from random, we can confidently conclude that this disease can be accurately predicted from microbiome data. It is noteworthy, however, that pitted scab disease is precisely one of the diseases that are easier to be visually detected, and thus, there is a reliable separation among the two classes (diseased and non-diseased) which is aiding in prediction by ML models. Other diseases, and more so yield, do not have such clear distinction between classes which results in lower predictive power. That is, we believe that the lack of prediction accuracy in yield is not driven by a lack of a biological connection between soil microbiome and yield, but on the lack of accurate labels that distinguish the two classes (e.g. low and high yield). If humans cannot distinguish what is low yield vs high yield, then that ambivalence will propagate into the ML classification. This conclusion seems to be confirmed when we notice that yield cannot be accurately predicted by any of 14 models in consideration (Table 2). We conclude that one of the main challenges when applying ML methods in biological applications is the artificial binarization of phenotypes. More work is needed to improve the performance of regression models that can predict continuous phenotypes when faced with limited sample sizes that are common in biological domains.
@@
We show that the normalization and zero replacement strategies for microbiome data sets have a huge impact on results, and unfortunately, we could not find any pattern to suggest that one strategy is superior to others. The FMS decision tree, however, suggests that certain normalization methods such as centered log-ratio normalization with pseudo-zero replacement (clr+pseudo) or centered log-ratio normalization with Bayesian-multiplicative treatment (clr+bayesMult) should not be used for disease prediction for future studies. We recommend domain scientists to perform different types of normalization and zero replacement strategies in their data prior to be used as input for prediction. See also [4, 22, 7] for other performance comparisons on types of normalizations.
@@
In terms of feature selection, we considered different strategies to select important features (ML, network comparison, and intersection of both). There is significant overlap between important OTUs by two methods and the inclusion of this subset of predictors allowed us to build less complex models with comparable good performance (see, for example, Figure S4 for pitted scab disease). We did not observe one feature selection strategy that outperformed the others. For example, the weighted F1 scores obtained when using the OTUs identified as important by the ML methods are comparable to the weighted F1 scores when including OTUs identified as important by network comparison. It is worth mentioning, however, that the performance is also comparable to that obtained when including all OTUs which shows that the feature selection strategies work at preserving the OTUs that have predictive signal while simultaneously allowing computationally expensive models (like Bayesian NN) to be applied. Tables S4 and S5 in the Supplementary Material contain the important OTUs identified by the two feature selection methods for the Phylum level associated with pitted scab disease.
@@
In our study, we could not identify an overall prediction superiority on certain taxonomic levels compared to others. While it is intuitive to expect that finer taxonomic levels would provide more power to predict potato outcomes, this is only true in some cases such as the OTU-S3 model in RF and the OTU-S1 model in Bayesian NN (Figure S4) where Genus level indeed outperforms the other taxonomic levels. However, we cannot find evidence to support this expectation in all other cases. Even though this finding disagrees with common expectations in microbiome research, we believe that it is a positive result from the ML perspective. If less refined taxonomic levels perform just as well as finer taxonomic levels, then the model could be fitted with fewer features, and would thus be less computational expensive and more efficient. Furthermore, the more features in the data, the more samples are needed to train the models. So, ML models that are trained on limited samples sizes (as the ones we have in this study) would benefit greatly by having fewer features.
@@
When including environmental features such as soil physicochemical properties and microbial population density of soil in the model, we achieved higher weighted F1 score values For Scabpit, utilizing alpha diversity with the RF model yields a median weighted F1 score of approximately 0.75. When combined with soil population density information, this score increases to 0.85. Similarly, for the Bayesian NN model, the score improves from 0.6 to 0.8 with the addition of soil population density information. For RF, employing OTU3 results in a median weighted F1 score of about 0.8, which further increases to 0.9 when supplemented with soil population density data. Similarly, for BNN, the median score reaches approximately 0.9. However, incorporating the population density of soil information leads to a narrower range for the box plot. In general, we investigated 14 different models for yield and disease prediction with different combinations of microbiome and environmental data. Results show poor performance in predicting yield across different models. The best results for pitted scab are achieved by combining alpha diversity and soil chemistry (Alpha+Soil) for RF and important OTUs and soil data (OTU-S3+Soil) for Bayesian NN. The median weighted F1 scores for predicting diseases range from 0.8 to 0.9 for RF and from 0.6 to 0.9 for Bayesian BNN models (refer to Figure 2). Although the best-performing models include microbiome predictors (Alpha and OTU-S3), it is important to note that the models without microbiome data are comparably powerful as those including microbiome data. Specifically, for the RF method, the F1 score exceeds 0.75 for Scabpit and Black_Scurf, and surpasses 0.6 for Scab and Scabsuper. While the models trained with microbiome data alone show that microbiome can effectively predict pitted scab disease, the fact that models without this type of data continue to perform well provides evidence that microbiome data may not be necessary to achieve reasonable prediction. In fact, when the collection of microbiome data requires much higher cost investment, the increased prediction accuracy by including microbiome predictors may not be enough to justify the extra cost.
@@
The RF and Bayesian NN methods offer similar performance in all models, which validates the results of both methods. However, in this study, Bayesian NN take far longer to train due to weight sampling and approximation, especially on more refined taxonomic level such as Order, Family, and Genus. Thus, we believe that based on the current dataset, RF is the preferred decision-making model with high prediction potential for disease outcomes when including a combination of microbiome and environmental predictors.
Methods
@@
Figure 1 shows a graphical representation of our pipeline with three major steps: (1) data preparation, (2) feature selection, and (3) classification based on random forest (RF) and Bayesian NN models with various types of predictors, including microbiome data (OTUs), environmental, and a combination of both. We describe each step in the pipeline in the next subsections.
Data description
@@
In this study, we focus on the soil microbiome (matrix of abundances) in a variety of taxonomic orders, including Phylum, Class, Order, Family, and Genus as well as other environmental information from soil samples acquired from potato fields in Wisconsin and Minnesota. The dataset consists of measurements related to soil health, potato yield and soil quality information. The soil health data are collected from pre-plant fields and include soil physicochemical properties, soil microbiome composition, soil microbiome diversity, and soil pathogen abundance. The potato yield and quality assessments are taken from the following growing season with indicators including tuber yield and disease severity. Overall, we have 256 samples, 108 of which are taken from fields in Minnesota, and 148 of which are taken from fields in Wisconsin. We list all measurements in Table S1 in the Supplementary Material.
Binarization of Response Variables
@@
All six responses in the dataset are continuous, so we need to binarize them to fit the classification models. For the disease-related responses, we simply make the binary label 00 if there is no presence of disease, and $1$ if there was detection of disease (that is, if the continuous response is greater than $0.0$). Binarizing the yield responses variables is harder as there is not a universal standard to classify potato yield to be low or high. Furthermore, yield values are highly dependent on the type of potato variety. We assign a label of 00 (low yield) to samples with a yield less than the variety-specific median. Similarly, we assign a label of $1$ (high yield) to samples with yield greater than the variety-specific median. We illustrate this approach in Figure S1 in the Supplementary Material.
@@
Data Filtering, Normalization, and Zero Replacement
The input data is a matrix with non-negative read counts that were generated by a sequencing procedure. Let $w^{(k)}=[w_{1}^{(k)},...,w_{p}^{(k)}]$ be the total read counts of sample $k$ containing $p$ OTUs, where $w^{(k)}$ is a composition that adds up to a fixed value of $m^{(k)}=\sum_{i=1}^{p}w_{i}^{(k)}$. This value $m^{(k)}$ is the sequencing depth, which varies across samples and is predetermined by technical factors resulting in highly sparse data. It is reasonable to filter out a certain set of OTUs as first data preparation step. For filtering, we only include OTUs that appear in at least 15 samples. Table S2 displays the number of features (OTUs) before and after filtering for different taxonomic levels.
@@
With four zero replacement methods and five normalization methods, we create 20 datasets by the combination of zero replacement and normalization techniques. Our goal is to study the effect of the zero replacement and normalization choice in the performance of the deep learning methods. Namely, we have the following 20 combinations, $\text{NM}_{1}$: TSS+none, $\text{NM}_{2}$: TSS+pseudo, $\text{NM}_{3}$: TSS+multRepl, $\text{NM}_{4}$: TSS+bayesMult, $\text{NM}_{5}$: CSS+none, $\text{NM}_{6}$: CSS+pseudo, $\text{NM}_{7}$: CSS+multRepl, $\text{NM}_{8}$: CSS+bayesMult, $\text{NM}_{9}$: COM+none, $\text{NM}_{10}$: COM+pseudo, $\text{NM}_{11}$: COM+multRepl, $\text{NM}_{12}$: COM+bayesMult, $\text{NM}_{13}$: rarefy+none, $\text{NM}_{14}$: rarefy+pseudo, $\text{NM}_{15}$: rarefy+multRepl, $\text{NM}_{16}$: rarefy+bayesMult, $\text{NM}_{17}$: clr+none, $\text{NM}_{18}$: clr+pseudo,, $\text{NM}_{19}$: clr+multRepl, and $\text{NM}_{20}$: clr+bayesMult. For convenience, we use the notation $\text{NM}_{i}$ (Normalization Method) for $i=1,\dots,20$ in the Full Model Selection section (See Section ‘Identifying general practices to predict potato disease from microbiome data using full model selection models’).
For the environmental predictors of soil chemistry and microbial population density in the soil, we apply six scaling methods: (1) standardize features by subtracting the mean and scaling to unit variance [8]; (2) scale each feature to a $[0,1]$ range; (3) scale each feature by its maximum absolute value; (4) scale features by subtracting the median and scaling to the interquantile range [32]; (5) transform the features to follow a normal distribution [21]; (6) normalize samples individually to the unit norm. After normalization, the datasets are split into training, validation, and testing sets with 10-fold cross-validation. We used 80% of samples for training and validation, and 20% for testing.
Data Augmentation
@@
Regarding the data augmentation procedure, instead of simply adding a randomly generated noise to the original sample, we subset the data by variety and label, compute the mean (and standard deviation) abundance value for this subset, and create a new sample that includes the original data plus a Gaussian error with mean $\mu/100$ and standard deviation $\sigma/100$ where $\mu,\sigma$ are the subset-specific mean and standard deviation, respectively. This approach is illustrated in Figure S2 in the Supplementary Material. By the end of this procedure we would have a balanced augmented training set with 400 samples per label for each of the five taxonomic levels (Phylum, Class, Order, Family, and Genus).
Feature Selection
Feature selection involves the identification of important features (or covariates) that have high predictive power. Given the high-dimensionality of the data (e.g. 256 original samples for 485 OTUs at the Genus level), feature selection is necessary, especially for Bayesian NN models that are computationally expensive.
We pursue two approaches for feature selection: (1) using ML models to assess variable importance, and (2) using network analyses. To focus exclusively on the effect of feature selection, we only consider one type of normalization and zero replacement strategy in this investigation, namely, total sum scaling normalization without zero replacement ($\text{NM}_{1}$: TSS+none).
Using ML Models for Feature Selection
@@
After running all seven strategies, we assign a value ("TOTAL") to each OTU based on the number of times the OTU is selected as an important feature under the seven criteria. That is, an OTU that is selected as important by all seven strategies will have a value of 7. The OTUs are sorted based on "TOTAL" column and top 30% of them are selected as important features. Thus, 30, 36, 75, 85, and 162 OTUs are selected for Phylum, Class, Order, Family and Genus levels, respectively. For example, Table S4 in the Supplementary Material shows the top 30 OTUs at Phylum level and by which strategies they are identified as important features to predict the pitted scab response.
Using Network Comparison for Feature Selection